OpenStudy (anonymous):

solve by the elimination method 0.3-0.2y=4 0.4+0.5y=-55/17 type an ordered pair simplify

OpenStudy (lalaly):

x is missing

jhonyy9 (jhonyy9):

0,3x -0,2y =4 / x 10 --- 3x -2y =40 0,4x +0,5y = -55/17 0,3x -0,2y =4 ------------- 0,1x +0,7y = -55/17 -4 1,7x +11,9y =-55 -68 1,7x +11,9y = -123 / x 10 17x + 119y = -1230 3x-2y =40 --- x=(40+2y)/3 17(40+2y)/3 +119y = -1230 680 +34y + 357y = - 3690 391y = - 4370 4370 y = - ------- 391

jhonyy9 (jhonyy9):

40 + 2(- 4370/391) 40-8740 - 8700 than x = ------------------ = ---------- = ------- 3 1173 1173

OpenStudy (anonymous):

no thats no right. simplify answer type an ordered pair